# Detecting and Filtering Unsafe Training Data via Data Attribution

### v1

@@ 3.2 Unsafe Training Data Filtering
Using the estimated influence scores, we perform retrieval to identify the training examples most influential on the unsafe target samples. We select the top $K$ elements from the ranked list of $\text{Inf}(z,\mathcal{D}_{\text{target}})$ values. Let $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})\subseteq\mathcal{D}_{\text{train}}$ denote the set of $K$ most influential training samples selected, when the target dataset $\mathcal{D}_{\text{target}}$ is used as the target for attribution. By removing $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ from the training data, we construct a cleaner dataset that is expected to improve model safety upon retraining.
@@ 4.1 Experimental Setup
Overview.
We focus on a realistic training scenario wherein the benign portion of the data can be heterogeneous, consisting of:
  1. 1.
    Fully benign prompt-response pairs with no harmful content.
  2. 2.
    Safe demonstrations: pairs where the prompt may be unsafe, but the response is aligned and refuses or mitigates the request. These “safe” demonstrations have been shown to improve model safety (Jain et al., 2023). We denote the set of safe demonstrations as $\mathcal{D}_{\text{safe}}$.
We augment this benign dataset with a small set of unsafe injections designed to induce harmful responses. Across all experiments, the total injection ratio is below $0.025\%$. Identifying these few harmful training instances is extremely challenging, but crucial for mitigating jailbreaking vulnerabilities.
Datasets
We use the dataset Ultrachat 200k as our benign dataset and consider two unsafe datasets, ToxicChat and XSTest-Response. For each unsafe dataset, we split the dataset into $\mathcal{D}_{\text{safe}}$ (unsafe prompts with safe responses), $\mathcal{D}_{\text{unsafe}}$ (unsafe prompts and unsafe responses), and $\mathcal{D}_{\text{test}}$ (held-out unsafe prompts for evaluation). In our experiments, we inject both the $\mathcal{D}_{\text{safe}}$ and $\mathcal{D}_{\text{unsafe}}$ from the unsafe dataset into the benign Ultrachat 200k dataset to form the whole training dataset $\mathcal{D}_{\text{train}}=\mathcal{D}_{\text{benign}}\cup\mathcal{D}_{\text{safe}}\cup\mathcal{D}_{\text{unsafe}}$.
  • Ultrachat 200k. Ultrachat 200k [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k is a heavily filtered version of the UltraChat Ding et al. (2023) dataset, which comprises over 200k instructions for instruction fine-tuning purposes. We use a subset of the train sft split for our $\mathcal{D}_{\text{benign}}$, comprising $41,573$ samples.
  • ToxicChat. ToxicChat [^3] https://huggingface.co/datasets/lmsys/toxic-chat Lin et al. (2023) is a dataset consisting prompt-response pairs annotated with prompt toxicity and jailbreak-ness (response toxicity), curated from user interactions. We use a subset of the latest version of ToxicChat: ToxicChat-0124, and apply the following split: $|\mathcal{D}_{\text{safe}}|=127$, $|\mathcal{D}_{\text{unsafe}}|=97$, and $|\mathcal{D}_{\text{test}}|=30$.
  • XSTest-Response. XSTest-Response [^4] https://huggingface.co/datasets/allenai/xstest-response Han et al. (2024) is a dataset consisting of responses to the original XSTest Röttger et al. (2023), which includes 446 annotations of prompt harmfulness and model response harmfulness. We use the official response_harmfulness subset and apply the following split: $|\mathcal{D}_{\text{safe}}|=121$, $|\mathcal{D}_{\text{unsafe}}|=65$, and $|\mathcal{D}_{\text{test}}|=20$.
@@ 4.2 Evaluation Metrics
Given the retrieved set $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ containing $K$ top influential training data to the validation set, we define the precision and recall as:
@@
We adopt the Area Under Precision Recall Curve (AUPRC) as well as the precision, recall and F1 scores calculated with the top 100 samples identified for a comprehensive evaluation of baselines models and our methods.
To evaluate model safety, we employ Attack Success Rate (ASR) on the test set $\mathcal{D}_{\text{test}}$, which measures the proportion of the unsafe prompts in $\mathcal{D}_{\text{test}}$ that successfully elicit unsafe responses from the model.
4.3 Baselines
@@
We include baselines from three categories: online API tools (OpenAI moderation API), fine-tuned LLM as detectors (Llama-Guard-3-8B, Wildguard) and other model-free methods (GradSafe).
OpenAI moderation.
@@
The OpenAI Moderation API Markov et al. (2023) is an online moderation tool that assess whether the content is unsafe across 11 safety genres. We take the binary prediction label from the model to calculate precision, recall and F1. For AUPRC we use the model’s highest confidence across all safety genres.
Llama-Guard-3-8B.
@@
Llama-Guard-3-8B Llama Team (2024) is a Llama-3.1-8B pretrained model, fine-tuned for content safety classification. For AUPRC we use the probability of outputting the token “unsafe”, consistent with previous methodologies Xie et al. (2024).
Wildguard.
@@
Wildguard Han et al. (2024) is an open one-stop moderation model trained on a Mistral-7B model that detects prompt harmfulness, response harmfulness, and whether the response is a refusal to the prompt. Similar to Llama-Guard-3-8B, we use the probability of outputting the token “unsafe” as the confidence to calculate AUPRC.
GradSafe.
@@
GradSafe Xie et al. (2024) differs fundamentally in methodology by analyzing gradients with respect to safety-critical parameters of Llama-2, specifically focusing on the gradient of the model’s compliance response to prompts. In contrast, our approach directly traces unsafe behaviors back to the training data by leveraging token-level attributions to identify the sources of unsafe outputs. GradSafe operates independently of the model’s responses, providing pre-hoc moderation akin to LLM classifiers, whereas our method emphasizes post-hoc attribution to uncover the origins of unsafe model behavior.
@@ 4.4 Results and Discussion
We first demonstrate that fine-tuning language models on jailbreaking training data can effectively compromise their safety. Table 1 presents the ASR across different models and datasets. In comparison to training on the benign dataset $\mathcal{D}_{\text{benign}}$ only, the ASR is significantly higher when training with $\mathcal{D}_{\text{train}}$ that consists of unsafe training data, confirming that training on injected unsafe data results in unsafe model behaviors.
@@
Model Data ToxicChat XSTest-Response
Llama-3-8B $\mathcal{D}_{\text{train}}$ 93.3% 50%
$\mathcal{D}_{\text{benign}}$ 66.7% 0%
Gemma-2-9B $\mathcal{D}_{\text{train}}$ 90.0% 100%
$\mathcal{D}_{\text{benign}}$ 83.3% 70%
Table 2: AUPRC of baseline models and our method. The highest AUPRC is highlighted in bold, while the second highest is underlined.
Method ToxicChat (%) XSTest-Response (%)
OpenAI Moderation API 11.7 11.8
Llama-Guard-3-8B 30.3 82.5
Wildguard 44.5 85.9
GradSafe 30.7 47.3
Llama-3-8B-DABUF 52.0 74.1
Gemma-2-9B-DABUF 49.1 64.1
Table 3: Precision, recall, and F1 scores of baseline models and our method, calculated based on the top 100 identified training data points. The highest F1 score is highlighted in bold, and the second highest is underlined.
ToxicChat XSTest-Response
Method Precision (%) Recall (%) F1 (%) Precision (%) Recall (%) F1 (%)
OpenAI Moderation API 16.0 16.5 16.2 19.0 28.8 22.9
Llama-Guard-3-8B 33.0 34.0 33.5 57.0 86.4 68.7
Wildguard 46.0 47.4 46.7 59.0 89.4 71.1
GradSafe 51.1 47.4 49.2 45.0 68.2 54.2
Llama-3-8B-DABUF 51.0 52.6 51.8 57.0 86.4 68.7
Gemma-2-9B-DABUF 52.0 53.6 52.8 53.0 80.3 63.9
Table 4: Attack Success Rate (ASR) comparison between models retrained with the top 100 unsafe training samples filtered by baseline methods and DABUF. A higher ASR reflects a more unsafe model.
Model Filtering Method ToxicChat XSTest-Response
Llama-3-8B OpenAI Moderation 96.7% 25%
GradSafe 93.3% 25%
Wildguard 90.0% 15%
Llama-Guard-3-8B 90.0% 50%
DABUF 86.7% 5%
Gemma-2-9B OpenAI Moderation 90% 45%
GradSafe 86.7% 60%
Wildguard 80% 20%
Llama-Guard-3-8B 90% 30%
DABUF 66.7% 15%
Table 2 and 3 respectively show the AUPRC and the top 100 precision, recall, and F1 of unsafe training data detection across models and methods. Our method demonstrates superior or comparable performance across different experimental settings. Notably, DABUF applied to Llama-3-8B achieves the highest performance in the ToxicChat injection experiment, showcasing the advantages of using data attribution for detecting unsafe training data. While state-of-the-art classifiers like Wildguard outperform DABUF in XSTest-Response experiments, this is largely due to XSTest-Response’s focus on toxicity and explicit harm, aligning with their in-distribution training data.
More importantly, Table 4 highlights that the proposed DABUF significantly outperforms all baseline methods in terms of the ASR of models retrained after filtering out the top 100 unsafe training samples identified by different methods. The results demonstrate that our data attribution approach, which explicitly accounts for the model training process, effectively identifies and filters unsafe training data that contributes the most to the unsafe model behaviors of interest, which leads to models with better safety when retrained on filtered datasets.
@@ 5.1 Problem Setup
Prior research (An et al., 2024) has shown that LLMs can exhibit gender biases, particularly in contexts such as hiring decisions. To explore this issue, we present a scenario where training data contains inherent gender biases. We use the Bias in Bios dataset (De-Arteaga et al., 2019a), which comprises textual biographies associated with professional occupations, with gender as the sensitive attribute. From the training split, we sampled a subset of $10,000$ biography-occupation pairs ($\mathcal{D}_{\text{benign}}$) and injected it with 150 biased biography-occupation pairs ($\mathcal{D}_{\text{unsafe}}$) to form the training set ($\mathcal{D}_{\text{train}}$). For the target set, we generated gender-biased model responses for 50 occupation prediction prompts ($\mathcal{D}_{\text{target}}$). Additional details of the experiment are provided in Appendix A. Since model responses in this scenario are relatively simple, we do not use externally identified unsafe training data as we did in jailbreaking setups.
@@ 5.2 Evaluation Metrics
Similar with jailbreaking injection, we evaluate the detection performance via precision, recall, F1 score, and AUPRC.
@@
where $TPR_{A=a}=p(\hat{Y}=1|A=a,Y=1)$ is the TPR for the gender group $A=a$. The TPR Gender Gap intuitively quantifies the extent to which a model’s predictions favor physician against nurse when the gender-related information in the prompt is switched from female to male. A larger TPR Gender Gap observed on the held-out test set indicates that the model exhibits stronger gender biases.
@@ 5.3 Results and Discussion
First, we demonstrate that training a model on gender-biased data leads to behaviors that reflect gender biases. As can be seen in Table 5, models trained on gender-biased data exhibit a significant higher TPR Gender Gap in comparison to those trained on unbiased data.
Table 5: TPR Gender Gap on unbiased and biased datasets.
Model Dataset TPR Gender Gap
Llama-3-8B Unbiased 0.04
Biased 0.16
Gemma-2-9B Unbiased 0.02
Biased 0.08
Table 6: AUPRC and precision, recall, and F1 of the baseline model and our methods. The highest AUPRC and F1 value are highlighted in bold, while the second highest underlined. The suffix DABUF denotes our method.
Method AUPRC Precision Recall F1
Ada3 0.089 0.500 0.330 0.400
Llama-3-8B-DABUF 0.474 0.610 0.407 0.488
Gemma-2-9B-DABUF 0.530 0.670 0.447 0.536
Table 7: TPR Gender Gap comparison between models retrained with the top 100 biased samples filtered by Ada3 and our method. A higher TPR Gender Gap indicates a model with more gender bias. The lowest value is highlighted in bold.
Model Filtering Method TPR Gender Gap
Llama-3-8B None 0.16
Ada3 0.28
DABUF 0.14
Gemma-2-9B None 0.08
Ada3 0.04
DABUF 0.02
Table 6 presents the AUPRC, precision, recall, and F1 score results. Embedding-based tools like Ada3 exhibit poor performance in detecting unsafe data, suggesting that general-purpose embedding tools are not well-suited for detecting gender-biased data. In contrast, as a model-free approach, our method adapts effectively to different injection scenarios using diverse unsafe validation data.
Table 7 presents the TPR Gender Gap for models retrained after removing the top 100 influential samples identified. These results demonstrate that our approach effectively detects training samples that contribute to gender bias, and retraining on the filtered data leads to models with reduced gender bias. Notably, while embedding-based methods like Ada3 achieve non-trivial detection performance, filtering using Ada3 actually results in a higher TPR Gender Gap compared to training on the original data for Llama-3-8B model. This suggests that embedding-based methods may not be suitable for training data filtering. We believe this occurs because embedding-based models focus primarily on semantics, causing them to remove both unsafe and safe training data indiscriminately. This indiscriminate removal undermines the model’s ability to effectively learn the nuances of the occupation prediction task.

### v2

@@ 3.4 Unsafe Training Data Filtering
Using the detection scores derived above, we filter the training dataset by removing training examples with highest scores. Specifically, for each $x\in\mathcal{D}_{\text{train}}$, we compute $s_{I^{*}}(x)$ and rank all training samples accordingly. We then select the top $K$ elements from this ranked list, denoting the selected subset as
$\mathcal{S}_{K}(\mathcal{D}_{\text{target}})=\arg\operatorname{top-k}_{x\in\mathcal{D}_{\text{train}}}\;s_{I^{*}}(x)$
By removing $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ from the training data, we obtain a cleaner dataset that is expected to reduce unsafe behaviors upon retraining.
@@ 4.1 Experimental Setup
Datasets.
We conduct our experiments following the standardized setup from DATE-LM Jiao et al. (2025)  [^3] https://huggingface.co/DataAttributionEval, which provides an evaluation setup for unsafe training data detection. This benchmark uses UltraChat-200k Ding et al. (2023) as the benign dataset, and injects unsafe data from ToxicChat Lin et al. (2023) and XSTest-Response Röttger et al. (2023) into UltraChat to form unsafe training data. ToxicChat and XSTest-Response cover a wide range of jailbreaking settings, including both toxicity/harmfulness and adversarial prompt-response pairs. In addition, we evaluate model safety using the official test splits. This ensures that our setting is directly comparable to prior baselines while maintaining consistency across datasets.
@@ 4.2 Evaluation Metrics
Detection.
Given the retrieved set $\mathcal{S}_{K}(\mathcal{D}_{\text{target}})$ containing $K$ top influential training data to the target set, we define the precision and recall as:
@@
We evaluate detection with the Area Under the Precision–Recall Curve (AUPRC).
Retraining.
To assess model safety before and after filtering, we report the Attack Success Rate (ASR) on the held-out test split $\mathcal{D}_{\text{test}}$, defined as the fraction of unsafe prompts that elicit an unsafe response. To obtain stable and reliable results, ASR is averaged over 10 independent inference runs on $\mathcal{D}_{\text{test}}$. The target set $\mathcal{D}_{\text{target}}$ is defined to be the union of all unsafe model outputs collected across these runs. For detailed evaluation protocol refer to Appendix A.
4.3 Methods for Comparison
Table 1: AUPRC ($\uparrow$) of baseline models and our proposed method. For each backbone model, the highest AUPRC among data attribution methods is highlighted in bold, while the second highest is underlined. The last column shows the average improvement ($\Delta$) of our methods over their respective baselines.
Model Method ToxicChat AUPRC (%) $\uparrow$ XSTest-Response AUPRC (%) $\uparrow$ $\Delta$
OpenAI Moderation API 24.3 37.8
Llama-Guard-3-8B 44.5 91.6
Wildguard 56.0 93.0
GradSafe 37.6 27.4
Llama-3-8B-Instruct GradSim 36.4 1.7
GradSim-DRA 83.2 81.5 +63.3
Repsim 56.3 88.5
Repsim-DRA 92.6 98.6 +23.2
Gemma-2-9B-it GradSim 41.1 50.2
GradSim-DRA 82.3 95.5 +43.3
Repsim 51.6 83.9
Repsim-DRA 98.2 99.8 +31.3
Baselines.
@@ 4.3 Methods for Comparison
We include baselines from four categories: online API tools (OpenAI moderation API), fine-tuned LLM as detectors (Llama-Guard-3-8B, Wildguard), model-free methods (GradSafe) and baseline data attribution methods (GradSim and RepSim).
@@
OpenAI moderation. The OpenAI Moderation API Markov et al. (2023) is an online moderation tool that assess whether the content is unsafe across 11 safety genres. We take the binary prediction label from the model to calculate precision, recall and F1. For AUPRC we use the model’s highest confidence across all safety genres.
@@
Llama-Guard-3-8B. Llama-Guard-3-8B Llama Team (2024) is a Llama-3.1-8B pretrained model, fine-tuned for content safety classification. For AUPRC we use the probability of outputting the token “unsafe”, consistent with previous methodologies Xie et al. (2024).
@@
Wildguard. Wildguard Han et al. (2024) is an open one-stop moderation model trained on a Mistral-7B model that detects prompt harmfulness, response harmfulness, and whether the response is a refusal to the prompt. Similar to Llama-Guard-3-8B, we use the probability of outputting the token “unsafe” as the confidence to calculate AUPRC.
@@
GradSafe. GradSafe Xie et al. (2024) detect unsafe training data by analyzing gradients with respect to safety-critical parameters of Llama-2, specifically focusing on the gradient of the model’s compliance response to prompts. GradSafe operates independently of the model’s responses, providing pre-hoc moderation akin to LLM classifiers, whereas our setting emphasizes post-hoc attribution to uncover the origins of unsafe model behavior.
GradSim. Previous work has demonstrated that GradSim Xia et al. (2024); Jiao et al. (2025) is effective in tracing the influence of training samples on a given target set. We denote $h(z)$ as the gradient of sample $z$. The Grad-Sim score is then defined as the inner product between the $\ell_{2}$-normalized $h(z)$ and the $\ell_{2}$-normalized target gradient mean $\mu_{Q}$.
RepSim. Representation Similarity Pezeshkpour et al. (2021) scores have been used for various tasks, such as training data selection and factual tracing Jiao et al. (2025). Here $h(z)$ is the last-layer hidden state of the final token in sample $z$. Similar with GradSim, the RepSim score is then defined as the inner product between the $\ell_{2}$-normalized $h(z)$ and the $\ell_{2}$-normalized target hidden state mean $\mu_{Q}$.
Our methods.
We apply the proposed DRA on GradSim and RepSim, respectively lead to GradSim-DRA and Repsim-DRA.
@@ 4.4 Results and Discussion
We first show the results of detection performance. Table 1 reports the AUPRC of unsafe training data detection across baseline classifiers and data attribution methods. Our proposed method consistently improves upon data attribution baselines that leverage either hidden states or gradients, often yielding substantial performance gains. Notably, after applying our approach, data-attribution-based detection not only surpasses its vanilla counterparts but also outperforms the strongest baseline classifiers, highlighting the effectiveness of data attribution for uncovering unsafe training signals that classifiers fail to capture.
Before we move on to the results of filtering and retraining, we demonstrate that fine-tuning language models on jailbreaking training data can effectively compromise their safety. Table 2 presents the ASR across different models and datasets. In comparison to training on the benign dataset $\mathcal{D}_{\text{benign}}$ only, the ASR is significantly higher when training with $\mathcal{D}_{\text{train}}$ that consists of unsafe training data.
@@
Model Data ToxicChat ASR $\downarrow$ XSTest-Response ASR $\downarrow$
Llama-3-8B-Instruct $\mathcal{D}_{\text{train}}$ 33.3% 73.0%
$\mathcal{D}_{\text{benign}}$ 26.7% 0.0%
Gemma-2-9B-it $\mathcal{D}_{\text{train}}$ 28.0% 71.0%
$\mathcal{D}_{\text{benign}}$ 15.7% 0.0%
Table 3 further shows that applying our method not only improves detection but also leads to safer retrained models, as shown by consistently lower ASR values. This underscores that our method is not only effective at identifying unsafe training data but also at filtering those samples that most strongly drive harmful model behaviors, leading to significant improvements in retrain safety. In particular, gradient-based attribution enhanced with our denoising and dimension selection approach yield the greatest increase in retrained model safety. We believe this is because gradient-based attributions can more accurately capture the underlying training dynamics.
Table 3: Attack Success Rate (ASR) comparison between models retrained with the top unsafe training samples filtered. A higher ASR reflects a more unsafe model and the lowest ASR for each setting is in bold. The last column shows the average decrease in ASR after applying DRA to baseline data attribution methods.
Model Filtering Method ToxicChat ASR $\downarrow$ XSTest-Response ASR $\downarrow$ $\Delta$
Llama-3-8B-Instruct Wildguard 29.0% 50.0%
GradSim 34.7% 68.0%
GradSim-DRA 26.0% 43.0% 16.9-16.9
Repsim 29.0% 63.0%
Repsim-DRA 27.3% 60.0% 2.35-2.35
Gemma-2-9B-it Wildguard 30.5% 30.0%
GradSim 29.0% 77.0%
GradSim-DRA 23.3% 3.0% 39.9-39.9
Repsim 29.0% 70.0%
Repsim-DRA 30.7% 64.0% 2.15-2.15
@@ 5.1 Problem Setup
Prior research (An et al., 2024) has shown that LLMs can exhibit gender biases, particularly in contexts such as hiring decisions. To explore this issue, we present a scenario where training data contains inherent gender biases. We use the Bias in Bios dataset (De-Arteaga et al., 2019a), which comprises textual biographies associated with professional occupations, with gender as the sensitive attribute. From the training split, we sampled a subset of $10,000$ biography-occupation pairs ($\mathcal{D}_{\text{benign}}$) and injected it with 150 biased biography-occupation pairs ($\mathcal{D}_{\text{unsafe}}$) to form the training set ($\mathcal{D}_{\text{train}}$). For the target set, we generated gender-biased model responses for 50 occupation prediction prompts ($\mathcal{D}_{\text{target}}$). Additional details of the experiment are provided in Appendix B.
@@ 5.2 Evaluation Metrics
Detection
Similar with jailbreaking injection, we evaluate the detection performance via AUPRC.
Retrain Safety
@@
where $TPR_{A=a}=p(\hat{Y}=1|A=a,Y=1)$ is the TPR for the gender group $A=a$. The TPR Gender Gap intuitively quantifies the extent to which a model’s predictions favor physician against nurse when the gender-related information in the prompt is switched from female to male. A larger TPR Gender Gap observed on the held-out test set indicates that the model exhibits stronger gender biases. Same with the jailbreaking setting, we average all results across 10 inference runs and the target set $\mathcal{D}_{\text{target}}$ is defined to be the union of all biased model outputs collected across these runs.
@@ 5.3 Results and Discussion
First, we demonstrate that training a model on gender-biased data leads to gender biases. As can be seen in Table 4, models trained on gender-biased data exhibit a significant higher TPR Gender Gap in comparison to those trained on unbiased data.
Table 4: TPR Gender Gap for model trained on unbiased and biased data.
Model Dataset TPR Gender Gap $\downarrow$
Llama-3-8B-Instruct Unbiased -0.066
Biased 0.848
Gemma-2-9B-it Unbiased 0.032
Biased 0.452
Table 5: Gender bias detection AUPRC. The highest AUPRC for each model is highlighted in bold, while the second highest is underlined. The last column reports the change ($\Delta$) compared to the corresponding baseline.
Model Method AUPRC $\uparrow$ $\Delta$
Llama-3-8B-Instruct Wildguard 0.059
GradSim 0.417
GradSim-DRA 0.550 +0.133+0.133
Repsim 0.333
Repsim-DRA 0.748 +0.415+0.415
Gemma-2-9B-it Wildguard 0.059
GradSim 0.404
GradSim-DRA 0.468 +0.064+0.064
Repsim 0.009
Repsim-DRA 0.369 +0.360+0.360
Table 6: TPR Gender Gap comparison between models retrained with the top 500 biased samples filtered. A higher TPR Gender Gap indicates a model with more gender bias. The lowest value for each model is highlighted in bold. The last column reports the change ($\Delta$) in gender gap when applying DRA to data attribution methods
Model Filtering Method TPR Gender Gap $\downarrow$ $\Delta$
Llama-3-8B-Instruct None 0.848
Wildguard 0.798
GradSim 0.650
GradSim-DRA 0.246 0.404-0.404
Repsim 0.296
Repsim-DRA 0.092 0.204-0.204
Gemma-2-9B-it None 0.452
Wildguard 0.480
GradSim 0.344
GradSim-DRA 0.224 0.120-0.120
Repsim 0.536
Repsim-DRA 0.346 0.190-0.190
Table 5 reports the AUPRC results for gender-biased unsafe data detection. We observe that moderation classifiers such as Wildguard fail to generalize to this setting, as gender bias often does not include explicit harm signals captured during moderator training. Applying our method consistently improves the performance of data attribution approaches, with particularly high gains for gradient-based method. This indicates that our approach is especially effective at amplifying the signal of subtle unsafe patterns like gender bias, where classifier-based tools struggle.
Table 6 presents the TPR Gender Gap for models retrained after removing the top 500 influential samples identified by different detection methods. These results show that applying our method to representation-based approaches consistently improves not only the detection of unsafe data but also the downstream model safety after retraining, even in subtle cases where unsafety is implicit, such as gender bias. In particular, gradient-based data attribution methods benefit the most from our approach: filtering guided by gradient-based attribution yields models with substantially reduced gender bias. This observation aligns with our findings in the jailbreak setting, further reinforcing that gradient-based attribution provides a principled solution for identifying and mitigating unsafe training data.